Given the following data, calculate product cost per unit under absorption costing.

Direct labor $ 13 per unit
Direct materials $ 7 per unit
Overhead
Total variable overhead $ 26,000
Total fixed overhead $ 96,000
Expected units to be produced 46,000 units

Multiple Choice

  • $20.00 per unit

  • $20.57 per unit

  • $22.09 per unit

  • $22.65 per unit

  • $24.00 per unit
Answer #1

variable overhead per unit=26,000/46,000=$0.57(Approx).

fixed overhead per unit=96,000/46,000=2.09(Approx).

Unit product cost=Direct materials+Direct labor+variable overhead+fixed overhead

=7+13+0.57+2.09

=$22.65 per unit.(Approx).